3-A car has wheels with a radius of 20cm. If the car is initially driving at 25 m/s, what angular acceleration would be required to stop the car within 75m? (a) −0.8 rad/s2 (b) −4.2 rad/s2 (c) −10.5 rad/s2 (d) −20.8 rad/s2

4. A force, under any circumstance, will produce a torque. (a) True (b) False

i got

3d

4b

is it correct
Answer #1

Ans -

3 (d)

4 (b)

Explanation :

3) v2 = u2 + 2as

As, v = 0

u = 25 m/s

s = 75 m

a = linear acceleration

On putting the values

(0)2 = (25)2 + 2a×75

a = - 25 / 6

a = - 4.17 m/s2

Angular acceleration is given by

= a / r

= - 4.17 / 0.2

= - 20.8 rad/s2

4) If the line of action of a force passes through the axis of rotation , it does not produce any torque.